### v1

# Sum of Squares Decompositions for Structured Biquadratic Forms

Abstract
This paper studies sum-of-squares (SOS) representations for structured biquadratic forms. We prove that diagonally dominated symmetric biquadratic tensors are always SOS. For the special case of symmetric biquadratic forms, we establish necessary and sufficient conditions for positive semi-definiteness of monic symmetric biquadratic forms, characterize the geometry of the corresponding psd cone, and prove that every such psd form is SOS. We also formulate conjectures regarding SOS representations for symmetric M-biquadratic tensors and symmetric B0-biquadratic tensors, discussing their likelihood and potential proof strategies. Our results advance the understanding of when positive semi-definiteness implies sum-of-squares decompositions for structured biquadratic forms.
Key words. Biquadratic forms, sum-of-squares, positive semi-definiteness, sos rank, bilinear forms.
AMS subject classifications. 11E25, 12D15, 14P10, 15A69, 90C23.
1 Introduction
A fundamental question at the intersection of algebra and optimization is whether a multivariate polynomial that is nonnegative everywhere (positive semi-definite, or psd) can be written as a sum of squares (sos) of polynomials. Hilbert [7] famously proved that this is not always true, except in a few special cases. This dichotomy between nonnegativity and a structured sos representation has profound implications, particularly because the existence of an sos decomposition makes the problem of verifying nonnegativity computationally tractable via semidefinite programming.
While the general problem remains open, significant research has focused on identifying broad classes of psd polynomials and tensors that are guaranteed to be sos. For instance, many structured tensors are known to be sos, including even-order weakly diagonally dominant tensors, symmetric M-tensors, and several other families mentioned in [12]. This paper focuses on a specific and important class of polynomials: biquadratic forms.
An $m\times n$ biquadratic form is a polynomial in two sets of variables, $\mathbf{x}\in\mathbb{R}^{m}$ and $\mathbf{y}\in\mathbb{R}^{n}$, that is quadratic in $\mathbf{x}$ when $\mathbf{y}$ is fixed, and quadratic in $\mathbf{y}$ when $\mathbf{x}$ is fixed. Such a form can be represented by a fourth-order tensor $\mathcal{A}=(a_{ijkl})$ as:
$P(\mathbf{x},\mathbf{y})=\sum_{i,k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}x_{i}x_{k}y_{j}y_{l}.$
A psd biquadratic form is one for which $P(\mathbf{x},\mathbf{y})\geq 0$ for all $\mathbf{x},\mathbf{y}$. It is sos if it can be written as a finite sum of squares of bilinear forms.
The question of whether every psd biquadratic form is sos was settled negatively by Choi [3], who provided an explicit $3\times 3$ counterexample. However, recent work [9] has shown that several important subclasses, such as weakly completely positive biquadratic tensors, are indeed sos. This positive result raises a natural and important question: What other structured psd biquadratic tensors are sos?
In particular, the authors of [8] recently identified and analyzed several new classes of psd biquadratic tensors, including diagonally dominated symmetric, symmetric M-, and symmetric $\mathrm{B}_{0}$-biquadratic tensors. A central question, previously unanswered, is whether the psd tensors in these classes also admit sos representations.
This paper provides a comprehensive study of this question. Our main contributions are:
  1. 1.
    We prove that every diagonally dominated symmetric biquadratic tensor is sos (Theorem 2.1).
  2. 2.
    We formulate conjectures on the sos property for symmetric M-biquadratic tensors (Conjecture 3.1) and symmetric $\mathrm{B}_{0}$-biquadratic tensors (Conjecture 4.1), discussing the likelihood of their resolution.
  3. 3.
    We focus on the special case of symmetric biquadratic forms. Notably, Choi’s counterexample [3] is not symmetric. We establish necessary and sufficient conditions for positive semi-definiteness of monic symmetric biquadratic forms in the $3\times 3$ case (Theorem 5.2), characterize the geometry of the corresponding psd cone (Theorem 5.3), and prove that every $3\times 3$ symmetric psd biquadratic form is SOS (Theorem 5.5). We further conjecture that such forms can be expressed as the sum of at most five squares of bilinear forms (Conjecture 5.6).
The rest of this paper is organized as follows. Section 2 is devoted to diagonally dominated tensors, where we prove Theorem 2.1. Sections 3 and 4 introduce M- and $\mathrm{B}_{0}$-biquadratic tensors and present our conjectures. Section 5 formally defines $x$-symmetric and $y$-symmetric biquadratic forms and studies monic symmetric biquadratic forms. Section 6 offers further remarks and discusses directions for future research.
We denote $(1,\cdots,m)$ as $[m]$.
2 Diagonally Dominated Symmetric Biquadratic Tensors
Suppose that $\mathcal{A}=(a_{ijkl})$, where $a_{ijkl}\in\mathds{R}$ for $i,k\in[m],j,l\in[n]$. Then $\mathcal{A}$ is called an $m\times n$ biquadratic tensor. If
$a_{ijkl}=a_{kjil}=a_{klij}$ (1)
for $i,k\in[m],j,l\in[n]$, then $\mathcal{A}$ is called a symmetric biquadratic tensor. Denote the set of all $m\times n$ biquadratic tensors by $BQ(m,n)$, and the set of all $m\times n$ symmetric biquadratic tensors by $SBQ(m,n)$. If the entries of $\mathcal{A}$ satisfy
$a_{ijij}\geq r_{ij}\equiv{1\over 2}\sum_{i_{2}=1}^{m}\left(\sum_{j_{2}=1}^{n}\left|\bar{a}_{iji_{2}j_{2}}\right|+\sum_{j_{1}=1}^{n}\left|\bar{a}_{ij_{1}i_{2}j}\right|\right),$ (2)
for all $i{\in[m]}$ and $j{\in[n]}$, then $\mathcal{A}$ is called a diagonally dominated biquadratic tensor. It was shown in [8] that a diagonally dominated symmetric biquadratic tensor is a psd tensor.
Theorem 2.1.
A diagonally dominated symmetric biquadratic tensor is an SOS tensor.
Proof.
Let $\mathcal{A}=(a_{ijkl})$ be a diagonally dominated symmetric biquadratic tensor. Consider the $mn\times mn$ matrix $M$ defined by
$M_{(i,j),(k,l)}=a_{ijkl},\qquad\text{for }i,k\in[m],\;j,l\in[n],$
where rows and columns are indexed by pairs $(i,j)\in[m]\times[n]$.
The symmetry conditions $a_{ijkl}=a_{kjil}=a_{klij}$ ensure that $M$ is symmetric. Indeed:
  • $M_{(i,j),(k,l)}=a_{ijkl}=a_{klij}=M_{(k,l),(i,j)}$ (symmetric)
  • $M_{(i,j),(k,l)}=a_{ijkl}=a_{kjil}$ (additional internal symmetry)
Now, define the vector $\mathbf{z}\in\mathds{R}^{mn}$ by $z_{(i,j)}=x_{i}y_{j}$ for $\mathbf{x}\in\mathds{R}^{m}$, $\mathbf{y}\in\mathds{R}^{n}$. Then the biquadratic form becomes:
$B(\mathbf{x},\mathbf{y})=\sum_{i,k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}x_{i}x_{k}y_{j}y_{l}=\mathbf{z}^{\top}M\mathbf{z}.$
We now analyze the diagonal dominance condition. For fixed $(i,j)$, the quantity $r_{ij}$ in (2) satisfies:
$r_{ij}$ $=\frac{1}{2}\sum_{i_{2}=1}^{m}\left(\sum_{j_{2}=1}^{n}|\bar{a}_{iji_{2}j_{2}}|+\sum_{j_{1}=1}^{n}|\bar{a}_{ij_{1}i_{2}j}|\right)$
$=\frac{1}{2}\left[\sum_{(i_{2},j_{2})\neq(i,j)}|M_{(i,j),(i_{2},j_{2})}|+\sum_{(i_{2},j_{1})\neq(i,j)}|M_{(i,j_{1}),(i_{2},j)}|\right].$
Since $M$ is symmetric and $M_{(i,j_{1}),(i_{2},j)}=M_{(i_{2},j),(i,j_{1})}$, the second sum equals the first sum. Therefore:
$r_{ij}=\sum_{(k,l)\neq(i,j)}|M_{(i,j),(k,l)}|.$
Thus, the diagonal dominance condition (2) becomes:
$M_{(i,j),(i,j)}\geq\sum_{(k,l)\neq(i,j)}|M_{(i,j),(k,l)}|\qquad\text{for all }(i,j)\in[m]\times[n],$
which means $M$ is a symmetric diagonally dominant matrix with nonnegative diagonal entries.
A classical result states that any symmetric diagonally dominant matrix with nonnegative diagonal entries is positive semidefinite. Hence $M\succeq 0$, and thus $B(\mathbf{x},\mathbf{y})=\mathbf{z}^{\top}M\mathbf{z}\geq 0$ for all $\mathbf{x},\mathbf{y}$, confirming that $\mathcal{A}$ is psd.
Moreover, such matrices admit an SOS decomposition. Specifically, any symmetric diagonally dominant matrix $M$ with nonnegative diagonals can be written as:
$M=\sum_{p=1}^{mn}\alpha_{p}\mathbf{e}_{p}\mathbf{e}_{p}^{\top}+\sum_{p<q}\beta_{pq}(\mathbf{e}_{p}+s_{pq}\mathbf{e}_{q})(\mathbf{e}_{p}+s_{pq}\mathbf{e}_{q})^{\top},$
where $\alpha_{p}\geq 0$, $\beta_{pq}\geq 0$, and $s_{pq}\in\{-1,+1\}$.
Now, for each term in this decomposition:
  • If $p=(i,j)$, then $\mathbf{e}_{p}\mathbf{e}_{p}^{\top}$ corresponds to $(x_{i}y_{j})^{2}$
  • If $p=(i,j)$ and $q=(k,l)$, then $(\mathbf{e}_{p}+s_{pq}\mathbf{e}_{q})(\mathbf{e}_{p}+s_{pq}\mathbf{e}_{q})^{\top}$ corresponds to $(x_{i}y_{j}+s_{pq}x_{k}y_{l})^{2}$
All these are squares of bilinear forms. Therefore:
$B(\mathbf{x},\mathbf{y})=\mathbf{z}^{\top}M\mathbf{z}=\sum_{r}({\bf w}_{r}^{\top}\mathbf{z})^{2},$
where each ${\bf w}_{r}^{\top}\mathbf{z}$ is a bilinear form in $\mathbf{x}$ and $\mathbf{y}$. This shows that $\mathcal{A}$ is an sos tensor. ∎
3 Symmetric M-Biquadratic Tensors
Let $\mathcal{A}=(a_{ijkl})\in SBQ(m,n)$, $\mathbf{x}\in\mathds{R}^{m}$ and $\mathbf{y}\in\mathds{R}^{n}$. Then $\mathcal{A}\cdot\mathbf{y}\mathbf{x}\mathbf{y})\in\mathds{R}^{m}$ and $(\mathcal{A}\cdot\mathbf{y}\mathbf{x}\mathbf{y}))_{i}=\sum_{k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}y_{j}x_{k}y_{l}$ for $i\in[m]$, $(\mathcal{A}\mathbf{x}\cdot\mathbf{x}\mathbf{y}))_{j}=\sum_{i,k=1}^{m}\sum_{l=1}^{n}a_{ijkl}x_{i}x_{k}y_{l}$ for $j\in[n]$. If there are $\lambda\in\mathds{R}$, $\mathbf{x}\in\mathds{R}^{m}$, $\|\mathbf{x}\|_{2}=1$, $\mathbf{y}\in\mathds{R}^{n}$, $\|\mathbf{y}\|_{2}=1$, such that
$\mathcal{A}\cdot\mathbf{y}\mathbf{x}\mathbf{y}=\lambda\mathbf{x},\ \mathcal{A}\mathbf{x}\cdot\mathbf{x}\mathbf{y}=\lambda\mathbf{y},$
then $\lambda$ is called an M-eigenvalue of $\mathcal{A}$, with $\mathbf{x}$ and $\mathbf{y}$ as its M-eigen vectors. M-eigenvalues of symmetric biquadratic tensors were introduced by Qi, Dai and Han [11] in 2009. It was proved there that a symmetric biquadratic tensor always has M-eigenvalues, and it is psd if and only if all of its M-eigenvalues are nonnegative.
Let $\mathcal{I}=(I_{ijkl})\in BQ(m,n)$, where
$I_{ijkl}=\left\{\begin{array}[]{cl}1,&\text{ if }i=k\text{ and }j=l,\\ 0,&\text{otherwise.}\end{array}\right$.
$\mathcal{I}$ is referred to as $M$-identity tensor in [5, 13].
Let $\mathcal{A}=(a_{ijkl})\in BQ(m,n)$. We call $a_{ijij}$ diagonal entries of $\mathcal{A}$ for $i{\in[m]}$ and $j{\in[n]}$. The other entries of $\mathcal{A}$ are called off-diagonal entries of $\mathcal{A}$.
A biquadratic tensor $\mathcal{A}$ in $BQ(m,n)$ is called a Z-biquadratic tensor if all of its off-diagonal entries are nonpositive. If $\mathcal{A}$ is a Z-biquadratic tensor, then it can be written as $\mathcal{A}=\alpha\mathcal{I}$-$\mathcal{B}$, where $\mathcal{I}$ is the M-identity tensor in $BQ(m,n)$, and $\mathcal{B}$ is a nonnegative biquadratic tensor. By [8], $\mathcal{B}$ has an M-eigenvalue. Denote ${\lambda_{\max}}(\mathcal{B})$ as the largest M eigenvalue of $\mathcal{B}$. If $\alpha\geq{\lambda_{\max}}(\mathcal{B})$, then $\mathcal{A}$ is called an M-biquadratic tensor. It was proved in [8], a symmetric biquadratic M-tensor is psd.
Conjecture 3.1.
A symmetric biquadratic M-tensor is an sos tensor.
This conjecture might be provable with substantial work, but would require developing new techniques to bridge the gap between M-eigenvalues and matrix eigenvalues.
4 Symmetric B0-Biquadratic Tensors
Let $\mathcal{A}=\left(a_{i_{1}j_{1}i_{2}j_{2}}\right)\in BQ(m,n)$. Suppose that for $i{\in[m]}$ and $j{\in[n]}$, we have
${1\over 2}\sum_{i_{2}=1}^{m}\left(\sum_{j_{2}=1}^{n}a_{iji_{2}j_{2}}+\sum_{j_{1}=1}^{n}a_{ij_{1}i_{2}j}\right)\geq 0,$ (3)
and for $i,i_{2}{\in[m]}$ and $j,j_{1},j_{2}{\in[n]}$, we have
${1\over 2mn}\sum_{i_{2}=1}^{m}\left(\sum_{j_{2}=1}^{n}a_{iji_{2}j_{2}}+\sum_{j_{1}=1}^{n}a_{ij_{1}i_{2}j}\right)\geq\max\left\{\bar{a}_{iji_{2}j_{2}},\bar{a}_{ij_{1}i_{2}j}\right\}.$ (4)
Then we say that $\mathcal{A}$ is a B0-biquadratic tensor. It was proved in [8] that a B0-biquadratic tensor is psd.
Conjecture 4.1.
A symmetric B0-biquadratic tensor is an sos tensor.
This conjecture is very unlikely to be provable with current knowledge; the conditions seem too weak to guarantee SOS decomposition.
For more discussion on biquadratic tensors, see [5, 13, 14]
5 Symmetric Biquadratic Forms
Suppose that $\mathcal{A}=(a_{ijkl})$, where $a_{ijkl}\in\mathds{R}$ for $i,k\in[m],j,l\in[n]$, is an $m\times n$ biquadratic tensor. Let
$P(\mathbf{x},\mathbf{y})=\sum_{i,k=1}^{m}\sum_{j,l=1}^{n}a_{ijkl}x_{i}y_{j}x_{k}y_{l},$ (5)
where $\mathbf{x}\in\mathds{R}^{m}$ and $\mathbf{y}\in\mathds{R}^{n}$. Then $P$ is called a biquadratic form. If $P(\mathbf{x},\mathbf{y})\geq 0$ for all $\mathbf{x}\in\mathds{R}^{m}$ and $\mathbf{y}\in\mathds{R}^{n}$, then $P$ is called positive semi-definite (psd). If
$P(\mathbf{x},\mathbf{y})=\sum_{p=1}^{r}f_{p}(\mathbf{x},\mathbf{y})^{2},$
where $f_{p}$ for $p\in[r]$ are bilinear forms, then we say that $P$ is sum-of-squares (SOS). The smallest $r$ is called the SOS rank of $P$. Clearly, $P$ is psd or SOS if and only if $\mathcal{A}$ is psd or SOS, respectively.
While a biquadratic form $P$ may be constructed from a biquadratic tensor $\mathcal{A}$, the tensor $\mathcal{A}$ is not unique to $P$. However, there is a unique symmetric biquadratic tensor $\mathcal{A}$ associated with $P$.
5.1 Historical Context and Symmetric Forms
In 1973, Calderón [2] proved that an $m\times 2$ psd biquadratic form can always be expressed as the sum of $\frac{3m(m+1)}{2}$ squares of bilinear forms. In 1975, Choi [3] gave a concrete example of a $3\times 3$ psd biquadratic form which is not SOS. Recently, in [4] it was shown that a $2\times 2$ psd biquadratic form can always be expressed as the sum of three squares, and in [10] that a $3\times 2$ psd biquadratic form can be expressed as the sum of four squares.
Motivated by the work of Goel, Kuhlmann, and Reznick [6] on even symmetric forms, we consider symmetric biquadratic forms.
Definition 5.1.
Suppose $P(\mathbf{x},\mathbf{y})=P(x_{1},\cdots,x_{m},y_{1},\cdots,y_{n})$. If
$P(x_{1},\cdots,x_{m},y_{1},\cdots,y_{n})=P(x_{\sigma(1)},\cdots,x_{\sigma(m)},y_{1},\cdots,y_{n})$
for any permutation $\sigma$, then $P$ is called $x$-symmetric. Similarly, we define $y$-symmetric biquadratic forms. If $P$ is both $x$-symmetric and $y$-symmetric, we call $P$ a symmetric biquadratic form.
Note that a symmetric biquadratic form $P$ has only four free coefficients. If we fix the diagonal coefficients as $1$, we may write such a form as
$P(\mathbf{x},\mathbf{y})=$ $\sum_{i=1}^{m}\sum_{j=1}^{n}x_{i}^{2}y_{j}^{2}+a\sum_{i\neq k}\sum_{j=1}^{n}x_{i}x_{k}y_{j}^{2}$
$+b\sum_{i=1}^{m}\sum_{j\neq l}x_{i}^{2}y_{j}y_{l}+c\sum_{i\neq k}\sum_{j\neq l}x_{i}y_{j}x_{k}y_{l}.$ (6)
We call such a form a monic symmetric biquadratic form. Note that if the diagonal coefficients are negative, then $P$ is not psd. If the diagonal coefficients are zero, then $P$ is not psd unless $a=b=c=0$. This exceptional cases are excluded from our analysis, as this paper focuses on psd and sos forms.
We call the unique symmetric biquadratic tensor corresponding to an $x$-symmetric, $y$-symmetric, symmetric, or monic symmetric biquadratic form a completely $x$-symmetric, completely $y$-symmetric, symmetric, or monic completely symmetric biquadratic tensor, respectively. Thus, we have several new subclasses of symmetric biquadratic tensors.
5.2 Positive Semi-definiteness of Monic Symmetric Forms
We now address the following question: Under which conditions on $a$, $b$, and $c$ is a monic symmetric biquadratic form $P$ psd?
Theorem 5.2.
Let $P$ defined by (6) be an $m\times n$ monic symmetric biquadratic form. Then $P$ is psd if and only if the parameters $a$, $b$, and $c$ satisfy:
  1. (i)
    $1-b+(m-1)(a-c)\geq 0$,
  2. (ii)
    $1-a+(n-1)(b-c)\geq 0$,
  3. (iii)
    $1-a-b+c\geq 0$ and $1-\frac{ab}{c}\geq 0$ if $c\neq 0$, $\frac{a}{c}\in[1-n,1]$, and $\frac{b}{c}\in[1-m,1]$,
  4. (iv)
    $1+(m-1)a+(n-1)b+(m-1)(n-1)c\geq 0$.
Proof.
Let $\mathcal{A}\in{SBQ(m,n)}$ be the biquadratic tensor corresponding to $P$. Then $\lambda$ is called an M-eigenvalue of $\mathcal{A}$, with $\mathbf{x}$ and $\mathbf{y}$ as its M-eigenvectors if $\|\mathbf{x}\|_{2}=1$, $\|\mathbf{y}\|_{2}=1$,
$\beta_{x}\mathbf{1}_{m}+(1-a+(b-c)(\alpha_{y}^{2}-1))\mathbf{x}=\lambda\mathbf{x},$
and
$\beta_{y}\mathbf{1}_{n}+(1-b+(a-c)(\alpha_{x}^{2}-1))\mathbf{y}=\lambda\mathbf{y}.$
Here, $\alpha_{x}=\mathbf{1}_{m}^{\top}\mathbf{x}$, $\alpha_{y}=\mathbf{1}_{n}^{\top}\mathbf{y}$, $\beta_{x}=\alpha_{x}(a+c(\alpha_{y}^{2}-1))$, $\beta_{y}=\alpha_{y}(b+c(\alpha_{x}^{2}-1))$, $\mathbf{1}_{m}$ and $\mathbf{1}_{n}$ denote the all ones vectors of length $m$ and $n$, respectively. Therefore, $\mathbf{x}$ is an M-eigenvector of $\mathcal{A}$ only if $\beta_{x}=0$ or $\mathbf{x}=\frac{1}{\sqrt{m}}\mathbf{1}_{m}$. Similarly, $\mathbf{y}$ is an M-eigenvector of $\mathcal{A}$ only if $\beta_{y}=0$ or $\mathbf{y}=\frac{1}{\sqrt{n}}\mathbf{1}_{n}$. In the following, we prove conditions (i)–(iv) by enumerating all M-eigenvectors.
For condition (i), take $\mathbf{x}=\frac{1}{\sqrt{m}}\mathbf{1}_{m}$ and let $\mathbf{y}$ satisfy $\beta_{y}=0$. Then we have $\lambda=1-b+(m-1)(a-c)$ is an M-eigenvalue of $\mathcal{A}$. Since $P$ is psd, it follows that $1-b+(m-1)(a-c)\geq 0$.
Condition (ii) follows by taking $\mathbf{y}=\frac{1}{\sqrt{n}}\mathbf{1}_{n}$ and letting $\mathbf{x}$ satisfy $\beta_{x}=0$. This leads to the M-eigenvalue $1-a+(n-1)(b-c)\geq 0$.
For condition (iii), let $\mathbf{x}$ satisfy $\beta_{x}=0$ and $\mathbf{y}$ satisfy $\beta_{y}=0$, respectively. A direct computation gives $\lambda=1-a-b+c\geq 0$ and $\lambda=1-\frac{ab}{c}\geq 0$ when $c\neq 0$, $\frac{a}{c}\in[1-n,1]$, and $\frac{b}{c}\in[1-m,1]$.
For condition (iv), let $\mathbf{x}=\frac{1}{\sqrt{m}}\mathbf{1}_{m}$ and $\mathbf{y}=\frac{1}{\sqrt{n}}\mathbf{1}_{n}$. A direct computation gives $\lambda=1+(m-1)a+(n-1)b+(m-1)(n-1)c\geq 0$.
Conditions (i)-(iv) have enumerated all M-eigenvalues. Based on the equivalence between positive semidefiniteness and the nonnegativity of the smallest M-eigenvalue, the conclusion of this theorem is obtained. This completes the proof. ∎
5.3 Geometry of the $3\times 3$ PSD Cone
Theorem 5.3.
Let $P$ be a $3\times 3$ monic symmetric biquadratic form as in Theorem 5.2. The set of coefficients $(a,b,c)$ for which $P$ is positive semi-definite is a convex polyhedron in $\mathbb{R}^{3}$ described by the four linear inequalities:
$1+2a+2b+4c\geq 0,$ (7)
$1+2a-b-2c\geq 0,$ (8)
$1+2b-a-2c\geq 0,$ (9)
$1-a-b+c\geq 0.$ (10)
The polyhedron has four faces, six edges, and four vertices. Its boundary consists of the planar faces where each inequality becomes an equality.
Proof.
Given that $c\neq 0$, $\frac{a}{c}\in[-2,1]$, $\frac{b}{c}\in[-2,1]$, and conditions (7)–(10) hold, it follows that $1-\frac{ab}{c}\geq 0$ is always true. Therefore, the psd condition is equivalent to the four linear inequalities (7)–(10). These define a convex polyhedron in $\mathbb{R}^{3}$. The faces are obtained by setting each inequality to equality while the others remain inequalities. The edges and vertices are intersections of these faces.
The explicit description of faces, edges, and vertices can be obtained by solving the linear system. For brevity, we omit the full enumeration here, but note that the polyhedron is bounded (unlike the outer approximation from the three conditions alone). ∎
5.4 Examples on the Boundary of the PSD Cone
As an illustration of Theorem 5.3, we present six explicit monic symmetric biquadratic forms, one on each edge of the psd cone. For each edge, we choose a simple interior point. The faces generated by conditions (7)–(10) are denoted by $F_{1},\dots,F_{4}$, respectively.
  1. 1.
    On edge $E_{12}$ (intersection of $F_{1}$ and $F_{2}$): $a=-\frac{1}{2}$, $b=0$, $c=0$:
    $P_{1}(\mathbf{x},\mathbf{y})=\sum_{i=1}^{3}\sum_{j=1}^{3}x_{i}^{2}y_{j}^{2}-\frac{1}{2}\sum_{i\neq k}\sum_{j=1}^{3}x_{i}x_{k}y_{j}^{2}.$
  2. 2.
    On edge $E_{13}$ (intersection of $F_{1}$ and $F_{3}$): $a=0$, $b=-\frac{1}{2}$, $c=0$:
    $P_{2}(\mathbf{x},\mathbf{y})=\sum_{i=1}^{3}\sum_{j=1}^{3}x_{i}^{2}y_{j}^{2}-\frac{1}{2}\sum_{i=1}^{3}\sum_{j\neq l}x_{i}^{2}y_{j}y_{l}.$
  3. 3.
    On edge $E_{14}$ (intersection of $F_{1}$ and $F_{4}$): $a=0$, $b=\frac{1}{2}$, $c=-\frac{1}{2}$:
    $P_{3}(\mathbf{x},\mathbf{y})=\sum_{i=1}^{3}\sum_{j=1}^{3}x_{i}^{2}y_{j}^{2}+\frac{1}{2}\sum_{i=1}^{3}\sum_{j\neq l}x_{i}^{2}y_{j}y_{l}-\frac{1}{2}\sum_{i\neq k}\sum_{j\neq l}x_{i}y_{j}x_{k}y_{l}.$
  4. 4.
    On edge $E_{23}$ (intersection of $F_{2}$ and $F_{3}$): $a=0$, $b=0$, $c=\frac{1}{2}$:
    $P_{4}(\mathbf{x},\mathbf{y})=\sum_{i=1}^{3}\sum_{j=1}^{3}x_{i}^{2}y_{j}^{2}+\frac{1}{2}\sum_{i\neq k}\sum_{j\neq l}x_{i}y_{j}x_{k}y_{l}.$
  5. 5.
    On edge $E_{24}$ (intersection of $F_{2}$ and $F_{4}$): $a=0$, $b=1$, $c=0$:
    $P_{5}(\mathbf{x},\mathbf{y})=(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})(y_{1}+y_{2}+y_{3})^{2}.$
  6. 6.
    On edge $E_{34}$ (intersection of $F_{3}$ and $F_{4}$): $a=1$, $b=0$, $c=0$:
    $P_{6}(\mathbf{x},\mathbf{y})=(x_{1}+x_{2}+x_{3})^{2}(y_{1}^{2}+y_{2}^{2}+y_{3}^{2}).$
These six forms are positive semi-definite and lie on the boundary of the psd cone (since they satisfy two equalities among the defining inequalities).
5.5 SOS Decompositions for Boundary Forms
We have explicit SOS representations for $P_{5}$ and $P_{6}$:
Proposition 5.4.
The forms $P_{5}$ and $P_{6}$ admit explicit representations as sums of squares of bilinear forms:
$P_{5}(\mathbf{x},\mathbf{y})$ $=(x_{1}y_{1}+x_{1}y_{2}+x_{1}y_{3})^{2}+(x_{2}y_{1}+x_{2}y_{2}+x_{2}y_{3})^{2}$
$\quad+(x_{3}y_{1}+x_{3}y_{2}+x_{3}y_{3})^{2},$
$P_{6}(\mathbf{x},\mathbf{y})$ $=(x_{1}y_{1}+x_{2}y_{1}+x_{3}y_{1})^{2}+(x_{1}y_{2}+x_{2}y_{2}+x_{3}y_{2})^{2}$
$\quad+(x_{1}y_{3}+x_{2}y_{3}+x_{3}y_{3})^{2}.$
Thus, $P_{5}$ and $P_{6}$ are SOS with sos rank at most three.
Numerical experiments via semidefinite programming confirm that $P_{1}$, $P_{2}$, $P_{3}$, and $P_{4}$ are also SOS. In particular, the form $P_{4}$ has been verified to be SOS, and its matrix representation $M$ has rank 5, suggesting an sos rank of at most 5.
Consider also the following example, which lies in the interior of the PSD cone:
$P_{0}(\mathbf{x},\mathbf{y})=(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})(y_{1}^{2}+y_{2}^{2}+y_{3}^{2}).$
This $3\times 3$ biquadratic form is a monic symmetric biquadratic form with $a=b=c=0$. It is psd and sos, and can be expressed as the sum of four squares of bilinear forms:
$P_{0}(\mathbf{x},\mathbf{y})=(x_{1}y_{1}+x_{2}y_{2}+x_{3}y_{3})^{2}+(x_{1}y_{2}-x_{2}y_{1})^{2}+(x_{1}y_{3}-x_{3}y_{1})^{2}+(x_{2}y_{3}-x_{3}y_{2})^{2}.$
This decomposition is minimal in the sense that the sos rank of $P_{0}$ is exactly four. Thus, any upper bound on the sos rank for the class of symmetric psd biquadratic forms must be at least four.
Theorem 5.5.
Every $3\times 3$ psd monic symmetric biquadratic form $P$ defined by (5) is a sum of squares of bilinear forms. Consequently, every $3\times 3$ psd symmetric biquadratic form is a sum of squares of bilinear forms.
Proof.
By Theorem 5.3, the set of coefficients $(a,b,c)$ for which the monic symmetric biquadratic form $P$ is PSD is a bounded convex polyhedron (a polytope) with vertices $V_{1}=\left(-\tfrac{1}{2},\ -\tfrac{1}{2},\ \tfrac{1}{4}\right)$, $V_{2}=\left(-\tfrac{1}{2},\ 1,\ -\tfrac{1}{2}\right)$, $V_{3}=\left(1,\ -\tfrac{1}{2},\ -\tfrac{1}{2}\right)$, and $V_{4}=\left(1,\ 1,\ 1\right).$ We first show that the forms corresponding to these vertices are SOS.
For $V_{4}$, we have $P_{V_{4}}(\mathbf{x},\mathbf{y})=\left(\sum_{i=1}^{3}x_{i}\right)^{2}\left(\sum_{j=1}^{3}y_{j}\right)^{2}$, which is the square of the bilinear form $(\sum_{i}x_{i})(\sum_{j}y_{j})$, hence SOS.
For $V_{1}$, we have the identity
$P_{V_{1}}(\mathbf{x},\mathbf{y})=\frac{1}{4}\left(\sum_{1\leq i<j\leq 3}(x_{i}-x_{j})^{2}\right)\left(\sum_{1\leq j<l\leq 3}(y_{j}-y_{l})^{2}\right),$
which can be expanded as
$P_{V_{1}}(\mathbf{x},\mathbf{y})=\frac{1}{4}\sum_{1\leq i<j\leq 3}\sum_{1\leq k<l\leq 3}\bigl((x_{i}-x_{j})(y_{k}-y_{l})\bigr)^{2},$
is a sum of squares of bilinear forms.
For $V_{2}$ and $V_{3}$, by symmetry it suffices to consider $V_{2}$. For $V_{2}$, we have
$P_{V_{2}}(\mathbf{x},\mathbf{y})=\frac{1}{2}\sum_{i=1}^{m}\left(x_{i}\sum_{j=1}^{n}y_{j}\right)^{2}+\frac{1}{2}\sum_{1\leq i<k\leq m}\left((x_{i}-x_{k})\sum_{j=1}^{n}y_{j}\right)^{2}$
is a sum of squares of bilinear forms. The same argument applies to $V_{3}$ by symmetry.
Thus all vertices of the PSD polytope are SOS. Since the polytope is the convex hull of its vertices and the set of SOS forms is a convex cone, every point in the polytope corresponds to an SOS form. This proves the first statement.
Now let $Q$ be an arbitrary $3\times 3$ symmetric biquadratic form that is PSD. By the symmetry, all its diagonal coefficients $q_{ijij}$ are equal. If they are all zero, then $Q$ is identically zero (as can be seen by evaluating on $x=e_{i},y=e_{j}$), hence SOS. Otherwise, we may scale $Q$ so that its diagonal coefficients become $1$. Then $Q$ becomes a monic symmetric biquadratic form, which by the first part is SOS. Scaling back preserves the SOS property. Therefore $Q$ is SOS. ∎
5.6 A Conjecture on $3\times 3$ Symmetric Forms
Based on these observations, we propose the following conjecture.
Conjecture 5.6.
A $3\times 3$ symmetric psd biquadratic form can always be expressed as the sum of at most five squares of bilinear forms.
Theorem 5.3 provides a complete characterization of the PSD polytope for $3\times 3$ monic symmetric forms, and Theorem 5.5 establishes that every such form is SOS. The explicit SOS decompositions presented here, along with the known sos rank of $P_{0}$ (four) and the numerical evidence for $P_{4}$ (five), suggest that the sos rank may be bounded by five. Resolving this conjecture would provide a precise quantitative understanding of the SOS complexity for this class.
6 Further Remarks
This work has advanced the understanding of the relationship between positive semi-definiteness and sum-of-squares representations for structured biquadratic forms. Theorem 2.1 establishes that diagonally dominated symmetric biquadratic tensors are not only psd but also sos. This provides a large and easily verifiable class of biquadratic forms for which nonnegativity is equivalent to an sos decomposition, making the property computationally tractable via semidefinite programming.
We have also introduced and formulated concrete conjectures regarding the sos property for two other important classes: symmetric M-biquadratic tensors (Conjecture 3.1) and symmetric B0-biquadratic tensors (Conjecture 4.1). We believe Conjecture 3.1 is likely true and may be provable by forging a deeper connection between the M-eigenvalue structure of these tensors and the spectral properties of their matrix unfoldings. In contrast, the weaker conditions defining B0-tensors seem insufficient to guarantee an sos decomposition with current techniques, making Conjecture 4.1 a more challenging and potentially false statement.
A particularly promising direction for future research lies in the study of symmetric biquadratic forms. By focusing on forms that are symmetric in the variables $\mathbf{x}$ and/or $\mathbf{y}$, we navigate away from known pathological counterexamples like Choi’s, which lacks such symmetry. The special case of $3\times 3$ monic symmetric biquadratic forms, fully characterized in Theorem 5.2 and whose PSD polytope is described in Theorem 5.3, serves as an ideal testbed. Theorem 5.5 establishes that every psd $3\times 3$ symmetric biquadratic form is sos, resolving the existence question for this class. Conjecture 5.6, which posits that every such form has sos rank at most five, is a natural and significant next target. A positive resolution would provide a precise quantitative understanding of the SOS complexity for this class, marking substantial advancement on a modern analogue of Hilbert’s 17th problem for a structured class of polynomials.
Remark on extending Theorem 5.5 to $m,n>3$. While Theorem 5.5 establishes that every psd $3\times 3$ symmetric biquadratic form is SOS, extending this result to general $m\times n$ dimensions presents significant challenges. For $3\times 3$, the PSD cone of monic symmetric forms is a convex polyhedron defined by four linear inequalities (Theorem 5.3), allowing a vertex-by-vertex verification that all extreme points are SOS. For $m,n>3$, however, the PSD conditions in Theorem 5.2 include nonlinear constraints (e.g., $1-ab/c\geq 0$ when $c\neq 0$), and the PSD set is no longer a polytope. Moreover, the M-eigenvalue characterization becomes more intricate, and explicit SOS constructions at the boundary of the PSD cone are not readily generalizable. Even if every psd symmetric form were SOS for all $m,n$, the minimal number of squares (the SOS rank) may grow with $m$ and $n$. Resolving the extension to higher dimensions remains an open problem, requiring new techniques from convex algebraic geometry and tensor spectral theory.
Acknowledgment We are thankful to Professor Giorgio Ottaviani, who proposed to study symmetric biquadratic forms. This work was partially supported by Research Center for Intelligent Operations Research, The Hong Kong Polytechnic University (4-ZZT8), the National Natural Science Foundation of China (Nos. 12471282 and 12131004), the R&D project of Pazhou Lab (Huangpu) (Grant no. 2023K0603), the Fundamental Research Funds for the Central Universities (Grant No. YWF-22-T-204), and Jiangsu Provincial Scientific Research Center of Applied Mathematics (Grant No. BK20233002).
Data availability No datasets were generated or analysed during the current study.
Conflict of interest The authors declare no conflict of interest.